a principal of $5000 is invested in an account paying an annual rate of 5%. find the amount after 6 years if the account is compounded semiannually, quarterly , and monthly.
Answer by DrBeeee(684)

Use your calculator to evaluate the following equations.
A) Compunded annually:
(1) B = 5000*(1 + 0.05)^(1*6) =
B) Compounded semi-annually:
(2) B = 5000*(1 + 0.05/2)^(2*6) =
C) Compounded quarterly:
(3) B = 5000*(1 + 0.05/4)^(4*6) =
D) Compounded monthly:
(4) B = 5000*(1 + 0.05/12)^(12*6) =
E) Compounded weekly:
(5) B = 5000*(1 + 0.05/52)^(52*6) =
I think you can see the trend don't you?